# Topological Big Bangs: Reflection, Itty-Bitty Blenders, and Eternal Trumpets

### v2

@@ Abstract
We discuss and formalize a topological means by which the initial singularity might be mollified, at the level of the spacetime manifold’s structure, in classical cosmological models of a homogeneous expanding universe. The construction, dubbed a “reflective” topological big bang, generalizes Schrödinger’s elliptic de Sitter space and is built to be compatible with the standard Friedmann-Lemaître-Robertson-Walker (FLRW) picture of the large-scale universe, only minimally modifying it via some nontrivial topology at an earliest “moment” in the universe’s history. We establish a mathematical characterization of the admissible topological structures of reflective topological big bangs, and we discuss implications for a standard concern in cosmology, the horizon problem. We present a nonreflective example that we’ve christened the Itty-Bitty Blender spacetime: this spacetime and its universal cover, the Eternal Trumpet spacetime, exhibit interesting potential structures of spacetimes avoiding the Hawking and Penrose singularity theorems.
@@ 1 Introduction
While the essential features of the big bang cosmological model, as realized in general relativity through the homogeoneous and isotropic Friedmann-Lemaître-Robertson-Walker (FLRW) spacetimes and perturbations thereof, have been remarkably successful in explaining a wide array of observable phenomena, physicists and cosmologists are broadly interested in avenues by which the initial singularity might be eliminated. Whatever the final theory of cosmology should be, physicists generally anticipate that the singularity will give way to something calculable. As quantum effects should increasingly dominate as one traces the universe’s expansion backwards, it is widely held that the singularity’s full resolution will be inextricably tied to a theory of quantum gravity. It may end up that the nature of this theory will render the classical picture of spacetime moot in this regime, but it is also possible that a manifold structure of spacetime continues to be meaningful, and one might then ask what kind of structure the manifold could have. A variety of possibilities for the nature of the early universe have been considered, with the most prominent propositions involving either inflation (very much the standard paradigm)– early references include [12, 20, 4, 11]; see [1] for a recent review of the theoretical and observational status of inflation– or (perhaps cyclic) “bouncing” cosmologies– reviews include [5, 8, 19]. See also [23, 30, 31]. In this work we present, at the classical level, a few alternative plausible structures of the spacetime manifold naturally compatible with FLRW metrics that morally amount to replacing the singularity with some nontrivial topology early in the universe’s evolution. Although they could in principle, these need not be thought of as supplanting an inflationary era. Indeed, as typical inflationary models are thought to be past-incomplete [7, 6, 13], they leave open the nature of the “initial” structure of the universe.
Figure 1: De Sitter space as a submanifold of Minkowski space in $n=2$. On the right, a pair of points to be identified is shown in a vertical cross-section.
@@
Under this foliation, the universe “bounces” from contraction to expansion at $t=0$, offering a schematic illustration of one possible mollification of the singularity. To obtain a qualitatively different mollification, one may modify de Sitter space by quotienting by antipodal identification, where antipodes $p\mapsto-p$ are taken in the ambient Minkowski space: this is elliptic de Sitter space, which we now denote by $M$. The contraction prior to the bounce and expansion thereafter are now identified, so that the hyperboloid is effectively reduced to a funnel shape, illustrated in Figure 2. The “mouth” of the funnel $M_{0}$ (the quotiented image of $\{0\}\times S^{n-1}$), acting as a sort of reflection surface, is now an $\mathbb{RP}^{n-1}$ instead of an $\mathbb{S}^{n-1}$, and the metric remains perfectly smooth across it since the antipodal identification was isometric. Due to this identification, a continuous curve which passes through the mouth apparently emerges on the opposite end of the spatial universe, as these opposite ends are smoothly joined along $M_{0}$.
@@
While elliptic de Sitter space is the best image to keep in mind, it has the perhaps less palatable pathology in $n=4$ (and indeed any even dimension) that it is nonorientable. In particular, this spacetime cannot admit continuous spinor fields [10], seemingly a most basic requirement if general relativity is to ultimately connect to quantum field theory. Although working reflective topological big bangs will always be closely associated with a symmetric bouncing cosmology (note that, due to the issue of orientability, the reverse is not true) obtainable by passing to a covering space, these provide qualitatively different pictures of the universe’s structure, so it is interesting to consider them separately and enumerate the possible ways of constructing a reflective topological big bang. In the following sections, we classify the possible orientable manifold structures compatible with a reflective topological big bang (Section 2), and we discuss these models’ relation to cosmology’s horizon problem as a means of illustrating their geometry (Section 3). In Section 4, we discuss an example, dubbed the “Itty-Bitty Blender” spacetime, of what one might call a topological big bang that is not reflective. This example is time-orientable, but it still admits closed timelike curves localized at the “beginning” of the universe. Its topologically trivial universal cover (the “Eternal Trumpet” spacetime), however, is globally hyperbolic. We discuss interplay with Hawking and Penrose’s celebrated singularity theorems [26, 15, 16, 17, 18] in cosmology, leading to several interesting natural questions as to whether the blender’s and trumpet’s structure can be replicated via more physical metrics. We summarize and conclude in Section 5.
2 Classifying Reflective Topological Big Bangs
@@
A smooth $n$-manifold $M$ is said to model a reflective topological big bang if it is diffeomorphic to $(\mathbb{R}\times\Sigma)/\mathbb{Z}_{2}$ for some connected (n1)(n-1)-manifold $\Sigma$, where the $\mathbb{Z}_{2}$ action is defined by $(t,p)\to(-t,\sigma(p))$ for some smooth free involution $\sigma:\Sigma\to\Sigma$.
@@
It is well-known that if a hypersurface is embedded in an orientable manifold, being one-sided is equivalent to being nonorientabile. As every nonorientable manifold admits a unique oriented double cover (the orientation double cover), this leads to our essential result, formalized below in Theorem 2.3, that the nonorientable reflection surface $M_{0}$ precisely and explicitly characterizes the topological structure of any orientable $M$ modeling a reflective topological big bang.
@@
Take $M=\bigwedge^{n-1}T^{*}\widetilde{\Sigma}$, and note that $\widetilde{\Sigma}$ naturally embeds into $M$ as the zero section $M_{0}$. Endowing $\widetilde{\Sigma}$ with some smooth Riemannian metric $\tilde{g}$, define $t:M\to\mathbb{R}$ by $t(\omega_{p})=\|\omega_{p}\|_{\tilde{g}}$, the $\tilde{g}-$norm of the top form $\omega_{p}$ based at $p\in\widetilde{\Sigma}$. We observe that $M_{+}:=t^{-1}(\mathbb{R}_{+})$ is an open submanifold. As each $t_{0}>0$ is a regular value of $t$, the level sets $M_{t_{0}}:=t^{-1}(t_{0})$ foliating $M_{+}$ are mutually diffeomorphic embedded submanifolds: denoting the hypersurface of unit top forms by $\Sigma:=M_{1}$, $M_{+}$ is apparently diffeomorphic to $\mathbb{R}_{+}\times\Sigma$.
@@
This theorem establishes that every possible orientable reflective topological big bang model may be uniquely and explicitly generated by considering the connected nonorientable manifolds $\widetilde{\Sigma}$, and the spatial topology $\Sigma$ of the corresponding expanding universe in each case is the orientation double cover of $\widetilde{\Sigma}$. Examples compatible with the possible homogeneous spatial geometries (constant positive, negative, or zero sectional curvature) may be constructed by seeking appropriate nonorientable quotients of $S^{n-1}$, $H^{n-1}$, and $\mathbb{R}^{n-1}$. It is interesting to note that in the case $n=4$ of greatest interest, and more generally for any even $n$, Synge’s theorem prohibits the existence of closed nonorientable (n1)(n-1)-manifolds with positive curvature, so there are no complete, oriented, constant positive curvature reflective topological big bang models (i.e. no working quotients of $S^{3}$).
@@
3 The Horizon Problem
In the previous section, we treated only the topological structure of a reflective topological big bang model. To showcase such a model’s geometric structure, we now discuss its interplay with cosmology’s horizon problem. This is a point on which a reflective model distinctly deviates from the related bouncing model.
@@ 3 The Horizon Problem
What might change in this picture in a reflective topological big bang? The most important point is that the spacetime manifold still perfectly admits the spatially homogeneous metric (3.1), with the only additional requirements for $C^{2}$ smoothness across the reflection surface being $\dot{a}(0)=0$ and $a_{0}:=a(0)>0$. According to equation (3.2), the condition $\dot{a}(0)=0$ requires that $\rho\to 0$ at the reflection surface. As we are working classically, we do not attempt to consider how this might occur in the complicated quantum-laden physics near the big bang for $a_{0}$ sufficiently small, but in a collection of classical perfect fluids, it would require at least one fluid to contribute with a negative energy density of magnitude comparable to the dominant positive component near $a_{0}$.
@@
with $\Omega_{ex}<<\Omega_{\gamma}$ to ensure agreement with standard cosmology except at the earliest times. Of course, we do not at all claim that equation (3.6) reflects a realistic model of the physical universe, but only that it is a minimal adjustment to standard cosmology yielding a geometry compatible with a reflective topological big bang, allowing analysis of qualitative features. That $p>4$ ensures that there is an earliest scale factor $a_{0}$ at which $\dot{a}=aH=0$. As this should occur long before matter-radiation equality, we neglect $\Omega_{m}$ and $\Omega_{\Lambda}$ to write
@@
Changing variables to $x:=\frac{\Omega_{ex}}{\Omega_{\gamma}}a^{-\epsilon}$, i.e. $a=\left(\frac{1}{x}\frac{\Omega_{ex}}{\Omega_{\gamma}}\right)^{1/\epsilon}=x^{-1/\epsilon}a_{0}$, we find
$\int_{0}^{t_{1}}\frac{dt}{a(t)}$ $\approx\frac{a_{0}}{H_{0}\sqrt{\Omega_{\gamma}}\epsilon}\int_{\frac{2}{2+\epsilon}}^{1}\frac{x^{-\frac{1}{\epsilon}}dx}{x\sqrt{1-x}}$
$>\frac{a_{0}}{H_{0}\sqrt{\Omega_{\gamma}}\epsilon}\int_{\frac{2}{2+\epsilon}}^{1}\frac{dx}{x\sqrt{1-x}}=\frac{a_{0}}{H_{0}\sqrt{\Omega_{\gamma}}}\cdot\frac{2}{\epsilon}\tanh^{-1}\left(\sqrt{\frac{\epsilon}{2+\epsilon}}\right).$
@@
$\frac{\Delta d_{causal}}{d_{CMB}}$ $\approx\frac{1}{d_{CMB}}\int_{0}^{t_{2}}\frac{dt}{a(t)}\approx\frac{a_{0}}{2\sqrt{a_{eq}}\epsilon}\int_{1/10}^{1}\frac{x^{-\frac{1}{\epsilon}}dx}{x\sqrt{1-x}}=\frac{a_{2}}{2\sqrt{a_{eq}}\epsilon}\int_{1/10}^{1}\frac{(10x)^{-\frac{1}{\epsilon}}dx}{x\sqrt{1-x}}$
$\leq\frac{a_{2}}{2\sqrt{a_{eq}}}\cdot\sup_{s>0}\left(s\int_{1/10}^{1}\frac{(10x)^{-s}dx}{x\sqrt{1-x}}\right)\approx 0.571\cdot\frac{a_{2}}{\sqrt{a_{eq}}}$
@@
meaning that the horizon problem cannot be resolved in the simple implementation of equation (3.6) without taking $a_{2}>a_{eq}$ and significantly impacting standard cosmology. This suggests that a reflective topological big bang cosmology still requires a process such as inflation in the early universe to resolve the horizon problem in a manner consistent with observations.

### v3

@@ Abstract
We discuss and formalize topological means by which the initial singularity might be mollified, at the level of the spacetime manifold’s structure, in classical cosmological models of a homogeneous expanding universe. One construction, dubbed a “reflective” topological big bang, generalizes Schrödinger’s elliptic de Sitter space and is built to be compatible with the standard Friedmann-Lemaître-Robertson-Walker (FLRW) picture of the large-scale universe, only minimally modifying it via some nontrivial topology at an earliest “moment” in the universe’s history. We establish a mathematical characterization of the admissible topological structures of reflective topological big bangs, and we discuss implications for a standard concern in cosmology, the horizon problem. We present a nonreflective example that we’ve christened the Itty-Bitty Blender spacetime: this spacetime and its universal cover, the Eternal Trumpet spacetime, exhibit interesting potential structures of spacetimes avoiding the Hawking and Penrose singularity theorems. While these toy models provide a proof-of-concept picture, several questions remain regarding the capacity to realize these structures under physical energy conditions.
@@ 1 Introduction
While the essential features of the big bang cosmological model, as realized in general relativity through the homogeoneous and isotropic Friedmann-Lemaître-Robertson-Walker (FLRW) spacetimes and perturbations thereof, have been remarkably successful in explaining a wide array of observable phenomena, physicists and cosmologists are broadly interested in avenues by which the initial singularity might be eliminated. Whatever the final theory of cosmology should be, physicists generally anticipate that the singularity will give way to something calculable. As quantum effects should increasingly dominate as one traces the universe’s expansion backwards, it is widely held that the singularity’s full resolution will be inextricably tied to a theory of quantum gravity. It may end up that the nature of this theory will render the classical picture of spacetime moot in this regime, but it is also possible that a manifold structure of spacetime continues to be meaningful, and one might then ask what kind of structure the manifold could have.
A variety of possibilities for the nature of the early universe have been considered, with the most prominent propositions involving either inflation (very much the standard paradigm)– early references include [13, 26, 4, 12]; see [1] for a recent review of the theoretical and observational status of inflation– or (perhaps cyclic) “bouncing” cosmologies– reviews include [5, 8, 21]. See also [30, 37, 38]. In this work we present, at the classical level, a few plausible structures of the spacetime manifold naturally compatible with FLRW metrics that morally amount to replacing the singularity with some nontrivial topology appended early in the universe’s evolution. Although they could in principle, these need not be thought of as supplanting inflation. Indeed, as typical inflationary models are thought to be past-incomplete [7, 6, 14], they leave open the nature of the “initial” structure of the universe. Even though the various models we present below are constructed to connect directly to FLRW geometries, then, we view them as propositions for plausible topological structures of the early universe that are independent of the existence of an inflationary era– such an era may or may not occur “after” these structures. That is, these are essentially variations on the FLRW framework within which one might try to implement reasonable physics, rather than specific instantiations of reasonable physics. Of course, if these frameworks are to convincingly operate as directly suggested by the given toy models, without a mechanism such as inflation, they must afford implementations of reasonable physics reproducing the various successes of inflation, such as resolving the horizon and flatness problems and generating a nearly scale-invariant power spectrum of primordial curvature perturbations. We do not claim to supply such implementations in this proof-of-concept work, and hence we also cannot currently speculate on unique observable signatures. We do, however, consider interplay of each model with the horizon problem, which is somewhat uniquely amenable to discussion at this level of breadth.
The remainder of this work is organized as follows. In Section 2, we introduce, define, and characterize “reflective” topological big bangs: We classify the possible orientable manifold structures compatible with them in Section 2.1, and we discuss their relation to the horizon problem as a means of illustrating their geometry in Section 2.2. These examples are not time-orientable, but this pathology is, in a sense, entirely localized to the “beginning” of the universe. In Section 3.1, we discuss an example, dubbed the Itty-Bitty Blender spacetime, of what one might call a topological big bang that is not reflective. This example is time-orientable, but it admits closed timelike curves also localized at the “beginning” of the universe. In Section 3.2, however, we discuss the Eternal Trumpet spacetime, the blender’s topologically trivial universal cover, which is globally hyperbolic. We summarize and conclude in Section 4.
Each spacetime explicitly written down in this work modeling a topological big bang violates standard energy conditions near the new structures appended to an FLRW spacetime. One may take the perspective that this is not particularly concerning since the actual matter in these regimes may well require a quantum mechanical description (perhaps depending on how early one appends the structures in question), rendering the pointwise imposition of energy conditions questionable. We consider this the best resolution of this concern for reflective topological big bangs. Precisely what makes the blender and trumpet examples interesting, however, is that their essential features do not necessitate the violation of physical energy conditions according to known theorems. Much of the discussion of Section 3.2 is devoted to these examples’ interplay with Hawking and Penrose’s celebrated singularity theorems [33, 17, 18, 19, 20] in cosmology, leading to several interesting open questions as to whether their structures can be replicated via more physically compelling metrics.
2 Reflective Topological Big Bangs
@@
Figure 1: De Sitter space as a submanifold of Minkowski space for $n=2$. On the right, a pair of points to be identified is shown in a vertical cross-section.
Under this foliation, the universe apparently “bounces” from contraction to expansion at $t=0$, offering a schematic illustration of one possible mollification of the singularity. To obtain a qualitatively different mollification, one may modify de Sitter space by quotienting by antipodal identification, where antipodes $p\mapsto-p$ are taken in the ambient Minkowski space: this is elliptic de Sitter space, which we now denote by $M$. The contraction prior to the bounce and expansion thereafter are now identified, so that the hyperboloid is effectively reduced to a funnel shape, illustrated in Figure 2. The “mouth” of the funnel $M_{0}$ (the quotiented image of $\{0\}\times S^{n-1}$), acting as a sort of reflection surface, is now an $\mathbb{RP}^{n-1}$ instead of an $\mathbb{S}^{n-1}$, and the metric remains perfectly smooth across it since the antipodal identification was isometric. Due to this identification, a continuous curve which passes through the mouth apparently emerges on the opposite end of the spatial universe, as these opposite ends are smoothly joined along $M_{0}$.
@@
While elliptic de Sitter space is the best image to keep in mind, it has the perhaps less palatable pathology in $n=4$ (and indeed any even dimension) that it is nonorientable. In particular, this spacetime cannot admit continuous spinor fields [11], seemingly a most basic requirement if general relativity is to ultimately connect to quantum field theory. Although working reflective topological big bangs will always be closely associated with a symmetric bouncing cosmology (note that, due to the issue of orientability, the reverse is not true) obtainable by passing to a covering space, these provide qualitatively different pictures of the universe’s structure, so it is interesting to consider them separately and enumerate the possible ways of constructing a reflective topological big bang.
2.1 Classifying Reflective Topological Big Bangs
@@
A smooth $n$-manifold $M$ is said to model a reflective topological big bang if it is diffeomorphic to $(\mathbb{R}\times\Sigma)/\mathbb{Z}_{2}$ for some connected (n1)(n-1)-manifold $\Sigma$, where the $\mathbb{Z}_{2}$ action is defined by $(t,p)\mapsto(-t,\sigma(p))$ for each $(t,p)\in\mathbb{R}\times\Sigma$, where $\sigma:\Sigma\to\Sigma$ is a smooth free involution.
@@
It is well-known that if a hypersurface is embedded in an orientable manifold, being one-sided is equivalent to being nonorientabile. As every nonorientable manifold admits a unique oriented double cover (the orientation double cover; see Theorem 15.42 of [25]), this leads to our essential result, formalized below in Theorem 2.3, that the nonorientable reflection surface $M_{0}$ precisely and explicitly characterizes the topological structure of any orientable $M$ modeling a reflective topological big bang.
@@
Take $M=\bigwedge^{n-1}T^{*}\widetilde{\Sigma}$, and note that $\widetilde{\Sigma}$ naturally embeds into $M$ as the zero section $M_{0}$. Endowing $\widetilde{\Sigma}$ with some smooth Riemannian metric $\tilde{g}$, define $t:M\to\mathbb{R}$ by $t(\omega_{p})=\|\omega_{p}\|_{\tilde{g}}$ for each top form $\omega_{p}\in M$ based at $p\in\widetilde{\Sigma}$. We observe that $M_{+}:=t^{-1}(\mathbb{R}_{+})$ is an open submanifold. As each $t_{0}>0$ is a regular value of $t$, the level sets $M_{t_{0}}:=t^{-1}(t_{0})$ foliating $M_{+}$ are mutually diffeomorphic embedded submanifolds: denoting the hypersurface of unit top forms by $\Sigma:=M_{1}$, $M_{+}$ is apparently diffeomorphic to $\mathbb{R}_{+}\times\Sigma$.
@@
This theorem establishes that every possible orientable reflective topological big bang model may be uniquely and explicitly generated by considering the connected nonorientable manifolds $\widetilde{\Sigma}$, and the spatial topology $\Sigma$ of the corresponding expanding universe in each case is the orientation double cover of $\widetilde{\Sigma}$. Examples compatible with the possible homogeneous spatial geometries (constant positive, negative, or zero sectional curvature) may be constructed by seeking appropriate nonorientable quotients of $S^{n-1}$, $H^{n-1}$, and $\mathbb{R}^{n-1}$. It is interesting to note that in the case $n=4$ of greatest interest, and more generally for any even $n$, Synge’s theorem (Corollary 3.10 in Section 9 of [10]) prohibits the existence of closed nonorientable (n1)(n-1)-manifolds with positive curvature, so there are no complete, oriented, constant positive curvature reflective topological big bang models (i.e. no working quotients of $S^{3}$).
@@
2.2 The Horizon Problem
So far, we have treated only the topological structure of a reflective topological big bang model. To showcase such a model’s geometric structure, we now discuss its interplay with cosmology’s horizon problem. This is a point on which a reflective model distinctly deviates from the related bouncing model.
@@ 2.2 The Horizon Problem
What might change in this picture in a reflective topological big bang? The most important point is that the spacetime manifold still perfectly admits the spatially homogeneous metric (2.2), with the only additional requirements for $C^{2}$ smoothness across the reflection surface being $\dot{a}(0)=0$ and $a_{0}:=a(0)>0$. According to equation (2.3), the condition $\dot{a}(0)=0$ requires that $\rho\to 0$ at the reflection surface. As we are working classically, we do not attempt to consider how this might occur in the complicated quantum-laden physics near the big bang for $a_{0}$ sufficiently small, but in a collection of classical perfect fluids, it would require at least one fluid to contribute with a negative energy density [^6] Bouncing cosmologies must also wrangle with the requirement $\dot{a}(0)=0$, and works in that direction have considered better-motivated matter models (see Section III.C of [5]), generally requiring a violation of the null energy condition near $t=0$. We invoke this simplest possible implementation to exhibit basic features and make a simple computation. of magnitude comparable to the dominant positive component near $a_{0}$.
@@
with $\Omega_{ex}<<\Omega_{\gamma}$ to ensure agreement with standard cosmology except at the earliest times. Of course, we do not at all claim that equation (2.7) reflects a realistic model of the physical universe, but only that it is a minimal adjustment to standard cosmology yielding a geometry compatible with a reflective topological big bang, allowing simple analysis of qualitative features. That $p>4$ ensures that there is an earliest scale factor $a_{0}$ at which $\dot{a}=aH=0$. As this should occur long before matter-radiation equality, we neglect $\Omega_{m}$ and $\Omega_{\Lambda}$ to write
@@
Changing variables to $u:=\frac{\Omega_{ex}}{\Omega_{\gamma}}a^{-\epsilon}$, i.e. $a=\left(\frac{1}{u}\frac{\Omega_{ex}}{\Omega_{\gamma}}\right)^{1/\epsilon}=u^{-1/\epsilon}a_{0}$, we find
$\int_{0}^{t_{1}}\frac{dt}{a(t)}$ $\approx\frac{a_{0}}{H_{0}\sqrt{\Omega_{\gamma}}\epsilon}\int_{\frac{2}{2+\epsilon}}^{1}\frac{u^{-\frac{1}{\epsilon}}du}{u\sqrt{1-u}}$
$>\frac{a_{0}}{H_{0}\sqrt{\Omega_{\gamma}}\epsilon}\int_{\frac{2}{2+\epsilon}}^{1}\frac{du}{u\sqrt{1-u}}=\frac{a_{0}}{H_{0}\sqrt{\Omega_{\gamma}}}\cdot\frac{2}{\epsilon}\tanh^{-1}\left(\sqrt{\frac{\epsilon}{2+\epsilon}}\right).$
@@
$\frac{\Delta d_{causal}}{d_{CMB}}$ $\approx\frac{1}{d_{CMB}}\int_{0}^{t_{2}}\frac{dt}{a(t)}\approx\frac{a_{0}}{2\sqrt{a_{eq}}\epsilon}\int_{1/10}^{1}\frac{u^{-\frac{1}{\epsilon}}du}{u\sqrt{1-u}}=\frac{a_{2}}{2\sqrt{a_{eq}}\epsilon}\int_{1/10}^{1}\frac{(10u)^{-\frac{1}{\epsilon}}du}{u\sqrt{1-u}}$
$\leq\frac{a_{2}}{2\sqrt{a_{eq}}}\cdot\sup_{s>0}\left(s\int_{1/10}^{1}\frac{(10u)^{-s}du}{u\sqrt{1-u}}\right)\approx 0.571\cdot\frac{a_{2}}{\sqrt{a_{eq}}}$
@@
meaning that the horizon problem cannot be resolved (at least in the simple implementation of equation (2.7)) without taking $a_{2}>a_{eq}$ and significantly impacting standard cosmology. This suggests that a reflective topological big bang cosmology still requires a process such as inflation in the early universe to resolve the horizon problem in a manner consistent with observations.